OpenStudy (anonymous):

Events A and B are mutually exclusive. Also, P(A) = 0.25 and P(B) = 0.35. What is P(A or B)? A.0.1 B.0.25 C.0.35 D.0.6 Please explain I don't get it

OpenStudy (ash2326):

Two events A and B are mutually exclusive this implies that if A occurs B can't occur and if B occurs A can't occur. So here P( A intersection B)=0 (intersection or simultaneous occurrence or and) We know that \[ P(A\ or\ B)= P(A)+P(B)- P( A\ and\ B)= 0.25+0.35-0=0.60\]

OpenStudy (anonymous):

ok thanks :)

OpenStudy (ash2326):

@Did you understand @Amy876 ?

OpenStudy (anonymous):

if they are mutually exclusive you can add the probabilities

OpenStudy (anonymous):

ya i got it thanks for ur help